OpenStudy (anonymous):

i need help with this problem... could someone help me with the steps please...Find a quadratic equation, with integer coefficients, that has solutions{3i,-3i} .

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow x^2 + 0x + 0 = -1 }\) The most basic form.

OpenStudy (anonymous):

it is a(x-3i)(x+3i) when a is any integer so the quardratic equa is a(x^2+9)

Parth (parthkohli):

You can add a more of bits. \(\Large \color{MidnightBlue}{\Rightarrow x^2 + 0x + 1 = 0 }\)

Parth (parthkohli):

Wait, its 3i..

Parth (parthkohli):

I guess you can manage the rest.

OpenStudy (anonymous):

it's wrong 'cuz -3i*3i has only one answer = -9i^2 = -9*-1 = 9

OpenStudy (anonymous):

i*i = -1

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

yes i remember that now :) thanks!!